OpenStudy (anonymous):

Solve for x. 3(x + 10) = -2(x - 15)

OpenStudy (anonymous):

3x+30=-2x+30 1x=0 x=0/1

OpenStudy (jwheele1):

1st distribute the three on the left then distribute the -2 on the right What do you get?

OpenStudy (anonymous):

@subbyh Thank you very much

OpenStudy (jwheele1):

3x + 30 = -2x + 30

OpenStudy (jwheele1):

isolate x on one side add 2x 5x +30 = 30

OpenStudy (anonymous):

\[3x+30=-2x+30\]

OpenStudy (jwheele1):

isolate x and its coefficient completely subtract 30 5x = 0

OpenStudy (anonymous):

i made a mistake it is not the answer i gave you 0/5

OpenStudy (anonymous):

i forgot to flip the sign

OpenStudy (jwheele1):

isolate x completely divide by 5 x = 0/5 OR 0

OpenStudy (anonymous):

@mermaidprincess13

OpenStudy (anonymous):

Ans: x=0.

OpenStudy (jwheele1):

Answer is 0

OpenStudy (anonymous):

sorry about that i did it in my mind but i just made sure on paper. what your main goal in an equation is to isolate variables on one side and numbers on the other to have like terms together which make calculations easier.

OpenStudy (anonymous):

i kinda noticed after i tried solving it myself. thanks for telling me. @subbyh